### v1

@@ 2.15 Novel Contribution: From Training-Time to Inference-Time Scaling
The originality of our work is that it establishes inference-time scaling laws as fundamental properties orthogonal to training-time scaling, and demonstrates that heterogeneous hardware orchestration can jointly optimize coverage, energy, latency, and cost through principled scaling law characterization. Our key finding—that 4.82× improvement in Intelligence Per Watt is achievable through principled scaling law optimization on heterogeneous hardware—demonstrates that the synthesis of these previously disjointed directions yields transformative efficiency gains, with simultaneously improved coverage (70% pass@k vs. 59.5%), energy reduction (47.7%), and latency improvement (22.5%) that exceed what any single technique can achieve. Critically, we address limitations identified in recent critiques of training-time scaling (Hooker, 2024) by implementing a practical, inference-time paradigm that operates without retraining, embraces hardware heterogeneity, and provides reliable predictive frameworks for deployment. This work thus positions inference-time scaling laws and heterogeneous orchestration as foundational concepts for sustainable edge AI deployment, offering a concrete technical realization of the theoretical future that recent scaling law critiques argue is necessary for continued progress.
@@ 3 Methodology
3.1 Foundational Framing: From Empiricism to Thermodynamics in AI
Before presenting QEIL’s technical framework, we establish a historical and conceptual analogy that motivates our approach. During the steam engine era, engineers observed that larger engines did more work and that higher boiler pressure increased output—empirical rules that powered the industrial revolution. These observations were captured as scaling laws: simple power-law relationships between pressure, size, and performance. Yet, for nearly a century, engineers operated these machines without understanding why these relationships held, what fundamental limits existed, or when scaling would fail. Boiler explosions, unexpected efficiency cliffs, and unpredictable failures plagued early industrial deployment because the underlying thermodynamic principles—entropy, energy conservation, irreversibility, and phase transitions—remained undiscovered.
@@ 3.1 Foundational Framing: From Empiricism to Thermodynamics in AI
The analogy to modern AI is direct and troubling. Today, we have empirical scaling laws for large language models: bigger models achieve lower loss, more data improves generalization, more compute enables emergent capabilities. These laws work remarkably well—we’ve built systems that genuinely scale and deliver value. Yet we operate at the steam engine era of AI: we lack deep theory explaining why scaling works, what fundamental capability limits exist, when scaling will break down, and how to guarantee safety at scale. On edge devices—where heterogeneous hardware introduces new constraints (memory bottlenecks, power ceilings, thermal throttling, latency hard caps)—these empirical laws often fail to predict outcomes or explain phase transitions in performance.
QEIL addresses this gap by introducing inference-time thermodynamic variables—analogs of entropy, temperature, potential, and efficiency limits—that extend classical scaling laws to edge environments. Just as thermodynamics provided engineers with predictive power, capability boundaries, and safety margins for steam engines, we propose that AI thermodynamics will provide researchers with:
  • Predictive depth: Understanding why inference scaling laws hold, rather than merely fitting curves to empirical data
  • Capability boundaries: Identifying fundamental limits on energy efficiency, latency, and accuracy—the Carnot limits of inference
  • Phase transition characterization: Explaining non-smooth behavior when quantization, thermal throttling, or memory bandwidth saturation occur
  • Edge deployment safety: Providing guarantees on when scaling is beneficial vs. dangerous in resource-constrained settings
This section presents QEIL as a step toward that theory. We do not claim QEIL is complete thermodynamics for AI—it is not. Rather, like Fourier’s law of heat conduction or Carnot’s cycle (which preceded a rigorous statistical mechanics foundation by decades), QEIL introduces thermodynamic-inspired frameworks that make edge constraints first-class citizens, enable precise optimization, and provide early warnings when systems approach limits.
@@
Our methodology unfolds in three integrated parts: (1) five theorems characterizing how coverage, energy, latency, and cost scale with fundamental parameters, proven empirically on real hardware and validated across model families; (2) energy-aware task decomposition, which decomposes inference into stages (embedding, decoder layers, output projection) aligned with their distinct hardware affinities and thermodynamic signatures; and (3) dynamic device orchestration, which uses greedy optimization to assign layers to heterogeneous devices, respecting memory, power, and latency constraints while minimizing dissipated energy.
@@ 3.2 System Architecture Overview
QEIL (Quantifying Edge Intelligence via Inference-time Scaling Laws) integrates inference-time scaling characterization with heterogeneous hardware orchestration to achieve unified optimization of coverage, energy, latency, and cost on edge devices. The framework combines three foundational components: (1) inference scaling law theorems characterizing how coverage and efficiency scale with model parameters, sample budget, and device characteristics, (2) energy-aware task decomposition breaking down inference into granular operations suitable for heterogeneous placement, and (3) dynamic device orchestration placing each task on the most cost-efficient hardware. This three-layer architecture enables systematic optimization of the total inference energy across heterogeneous CPU, GPU, and NPU devices while maintaining accuracy and meeting latency constraints.
@@ 3.2.1 Architecture Overview
The complete QEIL framework operates through three integrated stages as illustrated in Figure 1:
@@
Input Stage: The system accepts two primary inputs—model layer specifications including embedding dimensions, decoder layer configurations, and language model head parameters, as well as compute device specifications including maximum memory capacity ($M_{j}^{max}$), memory bandwidth ($B_{i}$), peak power consumption ($P_{i}$), compute frequency ($f_{i}$), device type (CPU, GPU, NPU), and priority ranking for load distribution.
@@
Energy-Aware Optimization Engine: This central component implements a multi-step optimization pipeline. First, preprocessing ranks available devices by energy efficiency, filtering devices that cannot accommodate the model constraints. Second, layer assignment logic strategically allocates the embedding layer and language model head to the most efficient device (typically CPU or NPU with high power efficiency). Third, decoder layers are distributed across remaining devices via greedy optimization that respects memory constraints while minimizing per-layer energy consumption. Finally, constraint checking validates that memory usage, latency SLA, and coverage targets are satisfied.
@@
Output Stage: The framework produces an optimal hardware-layer allocation list specifying which layers execute on which devices, along with predicted power consumption, efficiency factors (measured in accuracy per watt), estimated inference latency for prefill and decode phases, and maximum number of layers each device can accommodate given memory constraints. This allocation directly minimizes total energy expenditure ($\sum E_{stage_{i}}$) subject to device capacity and per-device performance constraints.
@@
Figure 1: QEIL (Quantifying Edge Intelligence via Inference-time Scaling Laws) Framework Architecture. Left panel shows model and device specifications as inputs. Center panel illustrates the three-stage optimization engine: (1) preprocessing and device ranking by efficiency, (2) layer assignment via greedy optimization with embedding/LM head selection and decoder layer distribution, and (3) constraint checking with helper functions computing power, efficiency, latency, and maximum layer capacity. Right panel outputs the optimal allocation plan. The objective function minimizes total inference energy across all heterogeneous devices.
3.2.2 Comparative Analysis: QEIL Information Density and Pass Rate Efficiency
A critical advantage of QEIL emerges when analyzing information density—the bits of information extracted per sample across different pass rate regimes. Figure 2 presents a comparative analysis of information encoding efficiency across four inference frameworks: (1) Pretraining—the baseline datacenter approach establishing a fixed information density of 0.5 bits/sample, (2) Inference Scaling (QEIL)—our edge-optimized framework extending baseline density to maintain consistent information extraction, (3) RL’s Information Density (Khatri et al., 2025)—reinforcement learning scaling that creates a peak density at moderate pass rates before collapsing, and (4) QEIL Enhanced Density—our optimized variant leveraging heterogeneous orchestration to maintain high information density across the entire pass rate spectrum.
The key insight is that information density reveals a fundamental trade-off invisible in raw accuracy metrics: most prior scaling approaches (RL training, pretraining) exhibit severe efficiency cliffs at extreme pass rates. At moderate pass rates ($10^{0}$$10^{1}$ %), QEIL maintains stable information density of approximately 0.5 bits/sample, while RL scaling peaks at 1.0 bits/sample at 15–20% pass rate, then catastrophically drops to 0.05 bits/sample at 100% pass rate. This represents a 20x efficiency collapse for RL methods, demonstrating that they sacrifice robustness at extreme confidence thresholds.
QEIL’s superiority stems from explicit optimization of information encoding across device boundaries:
  • Consistent information extraction: By decomposing inference into memory-bound (decode) and compute-bound (prefill) phases, QEIL ensures that each sample contributes uniform bits of information across all pass rate regimes, unlike RL training which concentrates information at intermediate pass rates
  • Pass rate robustness: QEIL Enhanced Density maintains 0.5–0.7 bits/sample even at extreme pass rates ($50\%$), while RL scaling collapses. This robustness is critical for safety-critical applications (medical diagnosis, autonomous systems) where the 99.99th percentile (near-100% accuracy) must be reliable
  • Heterogeneous hardware alignment: By placing low-information-density operations (embedding, sparse prefill) on NPUs and high-density operations (dense attention, KV cache decode) on GPUs, QEIL maximizes information per joule at each pass rate threshold
  • Sample diversity preservation: Classical scaling laws assume that each sample provides independent information. QEIL explicitly prevents information saturation by decomposing samples into information-dense stages, avoiding the peak-then-collapse pattern visible in RL training
The practical implication is stark: while pretraining provides baseline coverage with uniform information density, QEIL Enhanced Density achieves 40% higher information extraction at low pass rates ($1\%$), maintaining this advantage across the entire spectrum while avoiding RL’s catastrophic collapse at high pass rates. This validates that edge-aware inference optimization—combined with heterogeneous device placement—fundamentally outperforms datacenter-centric training approaches when information density (bits extracted per sample) is the optimization objective.
Figure 2: QEIL Information Density Across Pass Rates. The chart compares information extraction efficiency (bits of information per sample) as a function of pass rate (model confidence threshold) across four methods. Pretraining (baseline): Establishes flat information density at 0.5 bits/sample, the theoretical minimum for binary classification tasks. Inference Scaling (QEIL): Extends the baseline to higher pass rates, maintaining consistent density through heterogeneous orchestration. RL’s Information Density: Creates a peak at 15–20% pass rate (1.0 bits/sample) but suffers catastrophic collapse to 0.05 bits/sample at 100% pass rate, indicating severe overfitting to intermediate confidence thresholds and unreliability at extremes. QEIL Enhanced Density: Our proposed variant maintains high density (0.5–0.7 bits/sample) across the entire pass rate spectrum from 0.1% to 100%, preventing information saturation and ensuring robustness across all confidence regimes. The wide plateau of QEIL Enhanced Density (shown as dashed line) demonstrates that edge-aware inference scaling preserves information extraction efficiency even at extreme confidence thresholds, critical for safety-critical deployments.
This comparative framing—contrasting QEIL’s information-preserving inference scaling against classical training approaches—establishes the methodological foundation for the subsequent detailed theorems. The theorems formalize the information-theoretic variables, phase decomposition, and pass-rate-aware optimization principles that enable the information density gains demonstrated in Figure 2.
3.3 Inference Scaling Law Theorems
@@
The foundation of QEIL rests on five architecture-agnostic theorems characterizing how inference efficiency scales with fundamental parameters. These theorems are proven by empirical validation across models of varying parameter counts and device types, establishing universality independent of transformer architecture specifics.
3.3.1 Theorem 1: Coverage Scaling Law
@@
Theorem 1.1 (Inference Coverage Scaling). For any transformer-based language model with $N$ parameters generating $S$ samples of $T$ tokens per sample, the fraction of correctly solved queries (coverage) $C$ scales according to:
$C(S,N,T)=1-\exp\left(-\alpha(N)\cdot N^{\beta}\cdot S^{\beta}\cdot T^{\delta}\right)$ (1)
@@
where $\alpha(N)$ is a model-dependent coefficient (empirically $\alpha(N)\approx 10^{-4}$ for $N=125M$ to $20B$), $\beta\approx 0.7$ is the architecture-agnostic exponent (constant across all model families), $\delta\approx 0.2$ captures token length dependency, and $S$ is the number of samples.
Explanation: This law predicts that coverage improves log-linearly with sample budget, following a power-law relationship with model size. The exponentiated power law form emerges from information-theoretic principles: each additional sample provides diminishing marginal improvement (factor of $S^{\beta}$ with $\beta<1$) as common failure modes are exhausted. As the model size increases (exponent $\beta$ with respect to $N$), the baseline coverage improves. The temperature dependence $T$ reflects that longer outputs explore more reasoning paths, enabling higher coverage at the cost of increased latency and energy. Empirical validation across 1M queries on WILDCHAT and NATURALREASONING datasets confirms $\beta=0.7\pm 0.05$ and $\delta=0.2\pm 0.03$ hold universally across model families.
3.3.2 Theorem 2: Energy Scaling Law
@@
Theorem 2.1 (Inference Energy Scaling). For a language model of size $N$ (in parameters) executing inference on a device $i$ with peak power $P_{i}$ (watts), the total energy consumed across $S$ samples of $T$ tokens each scales as:
@@
  • $E_{0}(N)=c_{1}N^{\gamma}_{E}$ is the model-size-dependent base energy (per FLOP), with $\gamma_{E}\approx 0.9$ (sub-linear scaling due to improved cache locality at larger model sizes)
@@
Explanation: Energy scales linearly with sample count and token count because each token/sample incurs fixed computational cost (FLOPs for matrix multiplications). Energy scales sub-linearly with model size ($\gamma_{E}=0.9$) due to improved cache hit rates and memory reuse at larger model scales—the memory subsystem operates more efficiently. Quantization provides multiplicative reduction ($f(Q)$) by reducing data movement and improving cache efficiency. Device characteristics ($P_{i},\lambda_{i}$) reflect that heterogeneous devices have vastly different power envelopes and computational efficiency: a CPU consuming 45W can execute fewer FLOPS per joule than a specialized NPU consuming 5W. Hardware profiling across Intel NPU (60W TDP), NVIDIA GPU (300W), and CPU (45W) validates $\gamma_{E}=0.9$ and device multipliers empirically.
3.3.3 Theorem 3: Latency Scaling Law
Theorem 3.1 (Inference Latency Scaling). For sequential inference (single device) or parallel execution (heterogeneous orchestration), the end-to-end latency $\tau$ decomposes into distinct phases:

### v2

@@ 2.16 Novel Contribution: From Training-Time to Inference-Time Scaling
The originality of our work is that it empirically identifies inference-time scaling formalisms as fundamental properties orthogonal to training-time scaling, and demonstrates that heterogeneous hardware orchestration can jointly optimize coverage, energy, latency, and cost through principled scaling characterization. Our key finding—that 4.82× improvement in Intelligence Per Watt is achievable through principled energy optimization on heterogeneous hardware—suggests that the synthesis of these previously disjointed directions yields transformative efficiency gains, with simultaneously improved coverage (70% pass@k vs. 59.5%), energy reduction (47.7%), and latency improvement (22.5%) that exceed what any single technique can achieve. Critically, we address limitations identified in recent critiques of training-time scaling (Hooker, 2024) by implementing a practical, inference-time paradigm that operates without retraining, embraces hardware heterogeneity, and provides reliable predictive frameworks for deployment. Building on Asgar et al.’s foundational work on heterogeneous datacenter infrastructure, we extend these principles to resource-constrained edge environments, demonstrating that agentic orchestration across multi-vendor hardware enables practical edge AI deployment. Uniquely, we introduce a safety-first design paradigm that ensures reliable, fault-tolerant operation—bridging the gap between research prototypes and production-ready edge AI systems. This work thus positions inference-time scaling formalisms and heterogeneous orchestration as foundational concepts for sustainable, safe edge AI deployment, offering a concrete technical realization of the theoretical future that recent scaling critiques argue is necessary for continued progress.
@@ 3 Methodology
3.1 Foundational Framing: Empirical Scaling Formalisms for Edge AI
Before presenting QEIL’s technical framework, we establish the empirical and practical motivation for our approach. The history of engineering provides a useful analogy: during the steam engine era, engineers developed empirical rules relating pressure, size, and performance that powered the industrial revolution. These observations were captured as scaling relationships: power-law relationships that enabled practical engineering despite incomplete theoretical understanding.
@@ 3.1 Foundational Framing: Empirical Scaling Formalisms for Edge AI
The analogy to modern AI deployment is instructive. Today, we have empirical scaling relationships for large language models: bigger models achieve lower loss, more data improves generalization, more compute enables emergent capabilities. These relationships work remarkably well—we’ve built systems that genuinely scale and deliver value. However, on edge devices—where heterogeneous hardware introduces new constraints (memory bottlenecks, power ceilings, thermal throttling, latency hard caps)—these empirical relationships require careful validation and extension to predict outcomes in resource-constrained settings.
QEIL addresses this gap by introducing empirically-validated inference-time scaling formalisms that extend classical relationships to edge environments. Our approach provides:
  • Practical prediction: Empirically-grounded relationships that enable energy, latency, and coverage estimation for edge deployment planning
  • Operational boundaries: Clear characterization of the regimes where our scaling relationships hold and where they break down (e.g., under thermal throttling, memory pressure)
  • Safety margins: Explicit modeling of thermal constraints and reliability requirements that ensure safe operation in resource-constrained settings
  • Fault tolerance: Graceful degradation mechanisms when hardware constraints are exceeded or devices fail
@@
Our methodology unfolds in three integrated parts: (1) five scaling formalisms characterizing how coverage, energy, latency, and cost scale with fundamental parameters, validated empirically on real hardware across model families; (2) energy-aware task decomposition, which decomposes inference into stages (embedding, decoder layers, output projection) aligned with their distinct hardware affinities; and (3) safety-first agentic device orchestration, which uses greedy optimization to assign layers to heterogeneous devices, respecting memory, power, and thermal constraints while minimizing dissipated energy and ensuring reliable operation.
@@ 3.2 System Architecture Overview
QEIL (Quantifying Edge Intelligence via Inference-time Scaling Formalisms) integrates inference-time scaling characterization with heterogeneous hardware orchestration to achieve unified optimization of coverage, energy, latency, and cost on edge devices while maintaining system safety and reliability. The framework combines four foundational components: (1) inference scaling formalisms characterizing how coverage and efficiency scale with model parameters, sample budget, and device characteristics, (2) energy-aware task decomposition breaking down inference into granular operations suitable for heterogeneous placement, (3) dynamic device orchestration placing each task on the most cost-efficient hardware, and (4) safety and reliability monitoring that ensures thermal protection, fault tolerance, and adversarial robustness. This four-layer architecture enables systematic optimization of the total inference energy across heterogeneous CPU, GPU, and NPU devices while maintaining accuracy, meeting latency constraints, and guaranteeing safe operation.
@@ 3.2.1 Architecture Overview
The complete QEIL framework operates through four integrated stages as illustrated in Figure 1:
@@
Input Stage: The system accepts two primary inputs—model layer specifications including embedding dimensions, decoder layer configurations, and language model head parameters, as well as compute device specifications including maximum memory capacity ($M^{\max}_{j}$), memory bandwidth ($B_{i}$), peak power consumption ($P_{i}$), compute frequency ($f_{i}$), device type (CPU, GPU, NPU), thermal limits ($T^{\max}_{i}$), and priority ranking for load distribution.
@@
Energy-Aware Optimization Engine: This central component implements a multi-step optimization pipeline. First, preprocessing ranks available devices by energy efficiency, filtering devices that cannot accommodate the model constraints. Second, layer assignment logic strategically allocates the embedding layer and language model head to the most efficient device (typically CPU or NPU with high power efficiency). Third, decoder layers are distributed across remaining devices via greedy optimization that respects memory constraints while minimizing per-layer energy consumption. Finally, constraint checking validates that memory usage, latency SLA, coverage targets, and thermal safety margins are satisfied.
Safety and Reliability Monitor: This component continuously monitors device health and enforces safety constraints. It implements: (a) thermal throttling protection that reduces workload intensity when device temperatures exceed 80% of thermal limits; (b) fault detection and recovery that identifies device failures and redistributes workloads within 100ms; (c) input validation that rejects malformed or adversarial inputs; and (d) resource consumption bounds that prevent runaway inference from exhausting system resources. The safety monitor has override authority over the optimization engine—it can reject allocations that would compromise system safety, even if they are energy-optimal.
@@
Output Stage: The framework produces an optimal hardware-layer allocation list specifying which layers execute on which devices, along with predicted power consumption, efficiency factors (measured in accuracy per watt), estimated inference latency for prefill and decode phases, maximum number of layers each device can accommodate given memory constraints, and safety status indicators. This allocation directly minimizes total energy expenditure ($\sum E_{\text{stage}_{i}}$) subject to device capacity, per-device performance constraints, and safety requirements.
@@
Figure 1: QEIL (Quantifying Edge Intelligence via Inference-time Scaling Formalisms) Framework Architecture. Left panel shows model and device specifications as inputs. Center panel illustrates the four-stage optimization engine: (1) preprocessing and device ranking by efficiency, (2) layer assignment via greedy optimization with embedding/LM head selection and decoder layer distribution, (3) constraint checking with helper functions computing power, efficiency, latency, and maximum layer capacity, and (4) safety and reliability monitoring with thermal protection and fault tolerance. Right panel outputs the optimal allocation plan with safety guarantees. The objective function minimizes total inference energy across all heterogeneous devices subject to safety constraints.
3.3 Inference-Time Scaling Formalisms
@@
The foundation of QEIL rests on five empirically validated scaling formalisms characterizing how inference efficiency scales with fundamental parameters. These formalisms are derived from empirical validation across models of varying parameter counts and device types, establishing patterns within the transformer architecture family on our tested hardware platform. We emphasize that these are empirical relationships observed on our specific experimental setup; generalization to other architectures and hardware requires further validation.
3.3.1 Scaling Formalism 1: Coverage Scaling
@@
Formalism 1.1 (Inference Coverage Scaling). For transformer-based language models with $N$ parameters generating $S$ samples of $T$ tokens per sample, we observe that the fraction of correctly solved queries (coverage) $C$ scales according to:
$C(S,N,T)=1-\exp\left(-\alpha(N)\cdot N^{\beta_{N}}\cdot S^{\beta_{S}}\cdot T^{\delta}\right)$ (1)
@@
where $\alpha(N)$ is a model-dependent coefficient (empirically $\alpha(N)\approx 10^{-4}$ for $N=125M$ to 2.6B on our tested models), $\beta_{N}\approx 0.7$ is the model size scaling exponent, $\beta_{S}\approx 0.7$ is the sample count scaling exponent (we use separate exponents to allow independent characterization, though empirically they are approximately equal), $\delta\approx 0.2$ captures token length dependency, and $S$ is the number of samples.
Explanation: This formalism suggests that coverage improves with sample budget following a saturating exponential form. The exponentiated power law form captures the empirical observation that each additional sample provides diminishing marginal improvement (factor of $S^{\beta_{S}}$ with $\beta_{S}<1$) as common failure modes are exhausted. The separate treatment of model size ($\beta_{N}$) and sample count ($\beta_{S}$) allows more precise characterization, though we find empirically that $\beta_{N}\approx\beta_{S}\approx 0.7$ on our tested models. The token length dependence ($\delta\approx 0.2$) reflects that longer outputs explore more reasoning paths, with relatively modest impact compared to sample count—this aligns with the intuition that having more independent samples matters more than slightly longer individual samples. Empirical validation across our five model families indicates $\beta_{N}=0.70\pm 0.04$ and $\beta_{S}=0.70\pm 0.04$ with overlapping confidence intervals, suggesting these exponents are approximately equal within measurement precision.
Limitations: The confidence intervals for $\beta$ across models overlap substantially (see Table 1), meaning we cannot statistically distinguish model-specific exponents at our sample sizes. We interpret this as evidence that $\beta$ is approximately constant across tested models, not as proof of exact equality. Extrapolation beyond our tested range (125M–2.6B parameters) requires caution.
3.3.2 Scaling Formalism 2: Energy Scaling
@@
Formalism 2.1 (Inference Energy Scaling). For a language model of size $N$ (in parameters) executing inference on a device $i$ with peak power $P_{i}$ (watts), we find that the total energy consumed across $S$ samples of $T$ tokens each scales as:
@@
  • $E_{0}(N)=c_{1}N^{\gamma_{E}}$ is the model-size-dependent base energy (per FLOP), with $\gamma_{E}\approx 0.9$ (sub-linear scaling reflecting the empirical observation that larger models achieve better computational efficiency due to higher arithmetic intensity and reduced memory-bound overhead at larger batch sizes)
@@
Explanation: Energy scales linearly with sample count and token count because each token/sample incurs fixed computational cost (FLOPs for matrix multiplications). The sub-linear scaling with model size ($\gamma_{E}=0.9$) reflects that larger models have higher arithmetic intensity (more FLOPs per memory access) and benefit from better amortization of memory transfer overhead when batch sizes are appropriately scaled. Note: This is distinct from cache locality per se—larger models have larger working sets that may not fit in cache—but larger models operating at their optimal batch size achieve better compute utilization. Quantization provides multiplicative reduction ($f(Q)$) by reducing data movement and arithmetic precision. Device characteristics ($P_{i}$, $\lambda_{i}$) reflect that heterogeneous devices have vastly different power envelopes and computational efficiency. Hardware profiling across Intel NPU (25W TDP), NVIDIA GPU (300W TDP), and CPU (45W) validates the device multipliers empirically.
Energy Measurement Methodology: Energy measurements were obtained using a combination of hardware power monitoring and software instrumentation. For GPU measurements, we used NVIDIA’s nvidia-smi power queries sampled at 100ms intervals. For CPU and NPU measurements, we used Intel’s Running Average Power Limit (RAPL) interface accessed via the powercap sysfs interface. Total energy was computed by integrating instantaneous power over inference duration. All measurements were validated against external power meter readings (Watts Up Pro) with $5\%$ deviation.
3.3.3 Scaling Formalism 3: Latency Scaling
Formalism 3.1 (Inference Latency Scaling). For sequential inference (single device) or parallel execution (heterogeneous orchestration), we observe that the end-to-end latency $\tau$ decomposes into distinct phases: